OpenStudy (anonymous):

nine squared divided by three cubes. The given expression in symbols is and evaluate it.

OpenStudy (lgbasallote):

nine squared divided by 3 cubed is \[\huge \frac{9^2}{3^3}\] now can you solve it?

OpenStudy (lgbasallote):

let me make it easy for you \[\huge \frac{9 \times 9}{3\times 3 \times 3}\]

OpenStudy (anonymous):

Hint:\[\frac{9^2}{3^3}=\frac{(3^2)^2}{3^3}=\frac{3^4}{3^3}\]

OpenStudy (lgbasallote):

another hint \[\huge \frac{a^m}{a^n} = a^{m-n}\]